OpenStudy (anonymous):

problem help??? a. Find the slope of the line x + 4y = –8. b. Find the slope of the line that passes through the point (–4, -1) and is perpendicular to the line in part a. c. Write the equation of the line from part b in slope-intercept form.

OpenStudy (amistre64):

ax + by = c m = -a/b x +4y = 8; m = -1/4

OpenStudy (amistre64):

b is what???

OpenStudy (amistre64):

y = 4(x+4)-1; y = 4x +15

OpenStudy (amistre64):

thats a and c ...

OpenStudy (amistre64):

i spose b asks for the slope of the perp.... m = 4